# Minimal Fractional Topological Insulator in half-filled conjugate moiré Chern bands

### v2

@@ Abstract
We propose a “minimal” fractional topological insulator (mFTI), motivated by the recent experimental report on the signatures of FTI at total filling factor $\nu_{\rm tot}=3$ in a transition metal dichalcogenide moiré system. The observed FTI at $\nu_{\rm tot}=3$ is likely given by a topological state living in a pair of half-filled conjugate Chern bands with Chern numbers $C=\pm 1$ on top of another pair of fully-filled conjugate Chern bands. We propose the mFTI as a strong candidate topological state in the half-filled conjugate Chern bands. The mFTI is characterized by the following features: (1) It is a fully gapped topological order (TO) with 16 Abelian anyons if the electron is considered trivial (32 including electrons); (2) the minimally-charged anyon carries electric charge $e^{\ast}=e/2$, together with the fractional quantum spin-Hall conductivity, implying the robustness of the mFTI’s gapless edge state whenever time-reversal symmetry and charge conversation are present; (3) the mFTI is “minimal” in the sense that it has the smallest total quantum dimension (a metric for the TO’s complexity) within all the TOs that can potentially be realized at the same electron filling and with the same Hall transports; the mFTI is also the unique one that respects time-reversal symmetry. (4) the mFTI is the common descendant of multiple valley-decoupled “product TOs” with larger quantum dimensions. It can also be viewed as the result of gauging multiple symmetry-protected topological states. Similar mFTIs can be constructed for a pair of $1/q$-filled conjugate Chern bands. We classify the mFTIs via the stability of the gapless interfaces between them.
@@ I Introduction
The recent discovery of fractional Chern insulator Spanton et al. (2018); Xie et al. (2021) and fractional quantum anomalous Hall insulator at zero magnetic field Cai et al. (2023); Park et al. (2023); Zeng et al. (2023); Xu et al. (2023); Lu et al. (2023) have led to a new excitement on the strongly correlated states of matter in the moiré systems. Especially, the fractional topological insulator (FTI) reported most recently in transition metal dichalcogenide (TMD) homobilayer moiré heterostructure Kang et al. (2024), twisted bilayer MoTe2 to be more precise, represents a potentially new topological state of matter which could be fundamentally different from the previously known fractional quantum Hall state. In this TMD moiré system with spin-valley locking, the electron bands relevant to the experimental observations include two bands with Chern number $C=+1$ at valley-1 (spin-up), and two bands with opposite Chern number $C=-1$ at valley-2 (spin-down). Wu et al. (2019) The observed FTI signal appears at the total filling $\nu_{\rm tot}=3$. If the system preserves the time-reversal symmetry, and one ignores the fully occupied lower bands in each valley, the most relevant bands are a pair of time-reversal conjugate Chern bands with $C=\pm 1$, each with electron filling factor $\nu=1/2$. The experiment Kang et al. (2024) on twisted bilayer MoTe2 observed a series of appealing evidence that supports the existence of a certain type of time-reversal-invariant FTI with fractionalized helical edge states and spin-Hall conductivity $\sigma^{\rm sh}=1/2$ (when we only consider the contribution from the half-filled bands). The experimental evidence includes the nonlocal transport signal, the quantization of the edge conductance at fractional filling, and the suppression of the edge conductance when the time-reversal symmetry is explicitly broken by an in-plane magnetic field, analogous to the quantum spin Hall insulator König et al. (2007). It is worth noting that the relevant twisted bilayer MoTe2 system exhibits an $S_{z}$ spin quantum number conservation in addition to the charge conservation and time-reversal symmetry.Kang et al. (2024)
@@
The experimental facts have not yet completely pinned down the very nature of the reported FTI. A possiblity is that the FTI is simply a product of a pair of time-reversal conjugate topological orders (TO) in each valley, like some of the FTIs considered theoretically in the past Lindner et al. (2012); Cheng (2012) (For a review of previous theoretical discussion of FTI, please refer to Ref. Stern, 2016. We also refer to the seminal works Ref. Levin and Stern, 2009 and Levin and Stern, 2012 for general properties of generic FTIs). In the context of the TMD moiré system, the Chern band at valley-1 (valley-2) with Chern number $C=1$ ($C=-1$) can be viewed as a Landau level due to the real space pseudo magnetic field Morales-Dur¨¢n et al. (2023); Paul et al. (2023) (as illustrated in the upper panel of Fig. 1). For a half-filled Landau level, the candidate topological order of electrons could be the nonabelian Pfaffian state Moore and Read (1991); Greiter et al. (1991); Wen (1993), anti-Pfaffian state Levin et al. (2007), PH-Pfaffian state Son (2015), the Abelian $\mathrm{U}(1)_{8}$ state Greiter et al. (1992), or the “$331$” state Wen and Zee (1992); Overbosch and Wen (2008). Hence, the FTI could be the product of one of the TOs mentioned above in one valley, and its time-reversal conjugate in the other valley. In such product TOs, the TOs from the two valleys essentially decouple.
@@
In this work, we explore what kind of FTI is the “minimal” one, in the sense that it gives the desired experimental signals (especially the spin-Hall conductivity) and also has the minimal total quantum dimension (which is, in simple terms, a metric for the topological order’s complexity). We take charge conversation, spin-$S_{z}$ conservation, and time-reversal symmetry into account in the search for the minimal FTI. It turns out that the minimal FTI (mFTI) is unique, and it is NOT a simple product of a pair of conjugate TOs of electrons from the two valleys, though it can be viewed as the common descendant of multiple “product TOs” via anyon condensations. Our mFTI also has a spin-Hall conductivity $\sigma^{\mathrm{sh}}=1/2$, which is consistent with the experimental observation once the fully filled bands are included. This spin-Hall conductivity is also half of the value for the elementary non-interacting quantum spin-Hall insulators Kane and Mele (2005); Bernevig and Zhang (2006). The minimally charged anyon of the mFTI carries an electric charge $e^{\ast}=e/2$. In contrast, all the product TOs with the same response must have a minimal charge equal to or smaller than $e/4$. Additionally, according to the criterion established in Ref. Levin and Stern, 2009, our mFTI has robust gapless edge states that remain protected by time-reversal symmetry and charge conservation even when $S_{z}$ is not conserved.
@@ II Construction of the mFTI
The mFTI may be perceived and constructed in various ways. Let us first explore the most intuitive flux-attachment picture of the mFTI, which is based on the kinematics of both the inter-valley Cooper pair, and also the inter-valley exciton. In the TMD system, the motion of electrons in each valley is governed by a real-space pseudo magnetic field. Hence, the physics of a Chern band with Chern number $C=+1$ can be viewed as a Landau level in a magnetic field Morales-Dur¨¢n et al. (2023); Paul et al. (2023). More precisely, an electron in valley-1(2) will see magnetic flux ±$\pm\Phi_{0}=2\pi$ in each moiré unit cell. If one performs the standard flux attachment to electrons (independently for each valley), it is most natural to construct the conjugate composite Fermi liquid state discussed recently Zhang (2018); Myerson-Jain et al. (2023); Zhang (2024); Shi et al. (2024), which is a compressible state in the bulk. However, there is another alternative flux attachment which naturally leads to an incompressible FTI.
@@
indicating the desired spin-Hall conductivity $\sigma^{\mathrm{sh}}=1/2$. Also, there are no “diagonal” Hall responses for the electric charge and $S_{z}$ quantum numbers separately, which is consistent with time-reversal symmetry.
@@
The mutual Chern-Simons theory Eq. (1) implies that the mFTI is actually a $\mathbb{Z}_{4}$ TO “enriched” by two $\mathrm{U}(1)$ symmetries and time-reversal symmetry ${\cal T}$, i.e. it is a type of symmetry-enriched topological orders. Throughout this work, we assume that ${\cal T}^{2}=-1$ when acting on an electron or a hole, i.e. electrons/holes are Kramers doublets. Within the $\mathbb{Z}_{4}$ TO, the “minimally-charged” anyon $\mathbf{e}$ carries physical electric charge $e/2$, and the “minimal-spin” anyon $\mathbf{m}$ carries a physical spin quantum number $S_{z}=\hbar/4$. Both of these anyons are self-bosons, but they see each other as a $\Phi_{0}/4$ flux, i.e. they have mutual statistical angle $\pi/2$. The $\mathbf{e}\mathbf{m}$ bound state carries an electric charge and a spin quantum number $(e/2,\hbar/4)$, and it is an anyon with self-statistical angle $\pi/2$, i.e. its topological spin is $\theta_{\bf em}=e^{\mathrm{i}\pi/2}=\mathrm{i}$. This $\mathbb{Z}_{4}$ TO also preserves the time-reversal symmetry ${\cal T}$, which keeps the anyon $\mathbf{e}$ invariant and maps the anyon $\mathbf{m}$ to its anti-particle $\mathbf{m}^{3}$. Curiously, $\mathbf{m}^{2}$ is a Kramers doublet under time reversal with ${\cal T}^{2}=-1$. This result can be seen using both the edge theory in Sec. III and the general arguments of Ref. Levin and Stern, 2009, which we will elaborate more in Sec. IV.
@@
This $\mathbb{Z}_{4}$ TO have $16$ anyons modulo the electrons (and holes), which can be represented as composite states $\mathbf{e}^{p}\mathbf{m}^{q}$ with $p,q=0,\cdots 3$. However, since this $\mathbb{Z}_{4}$ TO is constructed with bosonic objects $b^{c}$ and $b^{s}$, the background gapped electrons and holes were not yet taken into account. Hence, there exist another 16 anyons, which are bound states of $\mathbf{e}^{p}\mathbf{m}^{q}$ and the electron. Note that the bound state of $\mathbf{e}^{p}\mathbf{m}^{q}$ and the electron has a different topological spin from $\mathbf{e}^{p}\mathbf{m}^{q}$. Here, we’ve effectively included the electron as a “transparent” Abelian anyon (that braids trivially with all other anyons), which is a convenient technical choice for showing the minimality of this $\mathbb{Z}_{4}$ TO later. With the electrons and holes included, the $\mathbb{Z}_{4}$ topological order has in total 32 Abelian anyons, resulting in a squared total quantum dimension ${\cal D}^{2}=32$. In Sec. IV, we will prove that ${\cal D}^{2}=32$ is the minimal squared total quantum dimension compatible with the given Hall transports associated with the electric charge and spin quantum number $S_{z}$. If one views electrons/holes as trivial anyons and identifies the anyons differed only by an electron/hole, then the mFTI effectively has a squared total quantum dimension $\tilde{{\cal D}}^{2}=16$. In the Appendix, we further prove that the mFTI proposed here is the unique mFTI compatible with charge conservation, $S_{z}$ conservation, time-reversal symmetry, and the desired Hall responses.
@@ III Experimental signatures
${\cal L}_{\mathrm{edge}}=\frac{4}{4\pi}\mathrm{i}\partial_{\tau}\phi^{c}\partial_{x}\phi^{s}-\frac{4}{4\pi}\mathrm{i}\partial_{\tau}\phi^{s}\partial_{x}\phi^{c}+\cdots,$ (3)
@@
If the spin $S_{z}$ conservation is broken but time-reversal symmetry is still preserved, an extra term $\cos(8\phi_{s})$ is allowed in the edge theory Eq. 3. When this term is relevant, it will lead to two-fold degenerate ground states at the edge, which are characterized by nonzero expectation values of gauge invariant and time-reversal odd operator $\cos(4\phi_{s})$, or $\sin(4\phi_{s})$. Hence, the edge theory remains gapless unless there is either spontaneous or explicit time-reversal symmetry breaking (assuming charge is conserved). A different way to understand the stability of the gapless edge using only the bulk data will be given below.
@@
${\cal L}_{\mathrm{edge}}=\frac{2}{4\pi}\mathrm{i}\partial_{\tau}\phi_{1}\partial_{x}\phi_{1}-\frac{2}{4\pi}\mathrm{i}\partial_{\tau}\phi_{2}\partial_{x}\phi_{2}-V_{IJ}\partial_{x}\phi_{I}\partial_{x}\phi_{J}.$ (4)
The density $\partial_{x}\phi_{1}$ (or $\partial_{x}\phi_{2}$) carries charge $e$ (or $-e$) and spin /2Planck-constant-over-$\hbar/2$. We have added the kinetic energy with the velocity matrix $V_{IJ}$ in Eq. (4) which arises from the non-topological part of the system. If $V_{IJ}$ is proportional to Pauli matrix $\sigma^{z}$, the left and right counter-propagating modes do not interact with each other, and this edge theory would lead to a fractionally quantized $1d$ conductance per edge:
@@
The smallest electric charge carried by the anyons of the mFTI is $e^{\ast}=e/2$. The odd value of $\sigma^{\rm sh}/e^{*}=1$ in the mFTI guarantees the gaplessness of the edge whenever time-reversal symmetry and charge conversation are present (even without the $S_{z}$ conservation).Levin and Stern (2009) In contrast, using similar arguments as Ref. Levin and Stern, 2009, one can show that, given the spin-Hall conductivity $\sigma^{\rm sh}=1/2$ (and the vanishing “diagonal” Hall responses associated with the charge and spin quantum numbers), any TO that is a product of two decoupled TOs from the two valleys must have excitations with fractional electric charges equal to or smaller than $e/4$. Therefore, the mFTI and the product TOs from the two valleys can be experimentally distinguished by probing the electric charge fractionalization. Examples of such probes include the shot noise of quantum point contactsSaminadayar et al. (1997, 1997), local capacitance probe of localized charge states Martin et al. (2004), and Coulomb oscillations in antidots Kou et al. (2012).

### v3

@@ Abstract
We propose a “minimal” fractional topological insulator (mFTI), motivated by the recent experimental report on the fractional quantum spin Hall effect in a transition metal dichalcogenide moiré system. The observed effect suggests the possibility of a topological state living in a pair of half-filled conjugate Chern bands with Chern numbers $C=\pm 1$. We propose the mFTI as a novel candidate topological state in the half-filled conjugate Chern bands. The mFTI is characterized by the following features: (1) It is a fully gapped topological order (TO) with 16 Abelian anyons if the electron is considered trivial (32 including electrons); (2) the minimally-charged anyon carries electric charge $e^{\ast}=e/2$, together with the fractional quantum spin-Hall conductance, implying the robustness of the mFTI’s gapless edge state whenever time-reversal symmetry and charge conservation are present; (3) the mFTI is “minimal” in the sense that it has the smallest total quantum dimension (a metric for the TO’s complexity) within all the TOs that can potentially be realized at the same electron filling and with the same Hall transports; the mFTI is also the unique minimal TO that respects time-reversal symmetry. (4) the mFTI is the common descendant of multiple valley-decoupled “product TOs” with larger quantum dimensions. It can also be viewed as the result of gauging multiple symmetry-protected topological states. Similar mFTIs are classified and constructed for a pair of $1/q$-filled conjugate Chern bands. We further classify the mFTIs via the stability of the gapless interfaces between them.
@@ I Introduction
The discovery of fractional Chern insulator Spanton et al. (2018); Xie et al. (2021) and fractional quantum anomalous Hall insulator at zero magnetic field Cai et al. (2023); Park et al. (2023); Zeng et al. (2023); Xu et al. (2023); Lu et al. (2023) have led to a new excitement on strongly correlated states of matter in the moiré systems. Recent breakthrough in twisted bilayer MoTe2 has found evidence of the fractional quantum spin Hall effect Kang et al. (2024), motivating the consideration of a potential fractional topological insulator (FTI), a new form of topological quantum matter which could be fundamentally different from the previously known fractional quantum Hall states. The electronic structure that underlies the fractional quantum spin Hall effect consists of conjugate Chern bands, namely time-reversed pairs of Chern bands with opposite Chern numbers.
Conjugate Chern bands are common in moiré systems based on transition metal dichalcogenide (TMD) homobilayers with spin-valley lockingWu et al. (2019); Devakul et al. (2021a); Crépel et al. (2024). A typical scenario contains a Chern band with Chern number $C=+1$ at valley-1 (spin-up) and its time-reversed partner, a Chern band with $C=-1$, at valley-2 (spin-down). In the twisted bilayer MoTe2 system experimentally studied in Ref. Kang et al., 2024, single and double integer quantum spin Hall insulators were observed at total fillings $\nu_{\rm tot}=2$ and $4$. These insulators naturally correspond to one and two pairs of fully filled conjugate Chern bands with $C=\pm 1$, consistent with the band structure calculationsJia et al. (2024); Wang et al. (2024); Xu et al. (2024). Note that the two pairs of conjugate Chern bands involved at $\nu_{\rm tot}=4$ share the same chirality within each valley. The fractional quantum spin Hall effect was discovered at $\nu_{\rm tot}=3$ where the first pair of conjugate Chern bands are fully filled and the second pair are half-filled. The observation includes nonlocal transport signals, the quantization of the edge conductance at fractional filling, and the suppression of the edge conductance when the time-reversal symmetry is explicitly broken by an in-plane magnetic field, analogous to the integer quantum spin Hall insulator König et al. (2007). These observed effects (modulo the contribution from the fully filled bands) motivate the consideration of a novel time-reversal invariant FTI with fractionalized helical edge states and spin-Hall conductance $\sigma^{\rm sh}=1/2$ in a pair half-filled conjugate Chern bands. In addition to fractionalization, the FTI is further embellished by the rich symmetries of relevant moiré systems, including charge conservation, spin-$S_{z}$ conservation, and time-reversal symmetryKang et al. (2024).
@@
The current experimental results in twisted bilayer MoTe2 are not yet sufficient to pin down the nature of the potential topological state. Also, conjugate Chern bands are ubiquitous in other moiré TMD systems as well. Hence, it is valuable to take a broader perspective and explore the landscape of possible FTI candidates in half-filled conjugate Chern bands. A possibility is that the FTI is simply a product of a pair of time-reversal conjugate topological orders (TO) in each valley, like some of the FTIs considered theoretically in the past Lindner et al. (2012); Cheng (2012) (For a review of previous theoretical discussion of FTI, please refer to Ref. Stern, 2016. We also refer to the seminal works Ref. Levin and Stern, 2009 and Levin and Stern, 2012 for general properties of generic FTIs). In the context of the TMD moiré system, the Chern band at valley-1 (valley-2) with Chern number $C=1$ ($C=-1$) can be viewed as a Landau level due to the real space pseudo magnetic field Morales-Dur¨¢n et al. (2023); Paul et al. (2023) (as illustrated in the upper panel of Fig. 1). For a half-filled Landau level, the candidate topological order of electrons could be the non-Abelian Pfaffian state Moore and Read (1991); Greiter et al. (1991); Wen (1993), anti-Pfaffian state Levin et al. (2007), PH-Pfaffian state Son (2015), the Abelian $\mathrm{U}(1)_{8}$ state Greiter et al. (1992), or the “$331$” state Wen and Zee (1992); Overbosch and Wen (2008). Some of such fractional quantum Hall states and the closely related composite Fermi liquids have been numerically investigated as the candidate states for the valley-polarized half-filled Chern bands in twisted bilayer MoTe2Xu et al. (2024); Ahn et al. (2024); Wang et al. (2024); Reddy et al. (2024); Chen et al. (2024); Dong et al. (2023); Goldman et al. (2023). Hence, the potential (valley-unpolarized) FTI in a pair of half-filled Chern bands with $C=\pm 1$ could be the product of one of the TOs mentioned above in one valley, and its time-reversal conjugate in the other valley. In such product TOs, the TOs from the two valleys essentially decouple. Understanding novel FTI candidates beyond such simple product TOs is both timely and important due to the potential relevance to the twisted bilayer MoTe2 and the ubiquity of conjugate Chern bands in moiré TMD systems. Also, the inter-valley interaction in the moiré TMD systems is expected to be strong, which often leads to valley polarization (which is necessary for the integer and fractional quantum anomalous Hall effects). In the system where the fractional quantum spin Hall effect was observed, although we do not expect valley polarization, the inter-valley interaction could still drive the system into novel FTIs fundamentally different from the product TOs of the two valleys, which would be the states when the intervalley interaction is ignored.
@@
In this work, we explore what kind of FTI in a pair of half-filled conjugate Chern bands with $C=\pm 1$ is the “minimal” one, in the sense that it gives the desired experimental signals (especially the spin-Hall conductance) and also has the minimal total quantum dimension (which is, in simple terms, a metric for the topological order’s complexity). We take charge conservation, spin-$S_{z}$ conservation, and time-reversal symmetry into account in the search for the minimal FTI. It turns out that the minimal FTI (mFTI) is unique, and it is NOT a simple product of a pair of conjugate TOs of electrons from the two valleys, though it can be viewed as the common descendant of multiple “product TOs” via anyon condensations. Our mFTI also has a spin-Hall conductance $\sigma^{\mathrm{sh}}=1/2$, which is consistent with the experimental observation once the fully filled bands are included. This spin-Hall conductance is also half of the value for the elementary non-interacting quantum spin-Hall insulators Kane and Mele (2005); Bernevig and Zhang (2006). The minimally charged anyon of the mFTI carries an electric charge $e^{\ast}=e/2$. In contrast, all the product TOs with the same response must have a minimal charge equal to or smaller than $e/4$. Additionally, according to the criterion established in Ref. Levin and Stern, 2009, our mFTI has robust gapless edge states that remain protected by time-reversal symmetry and charge conservation even when $S_{z}$ is not conserved.
@@ II Construction of the mFTI
The mFTI may be perceived and constructed in various ways. Let us first explore the most intuitive flux-attachment picture of the mFTI, which is based on the kinematics of both the inter-valley Cooper pair, and also the inter-valley exciton. In the TMD system, the motion of electrons in each valley is governed by a real-space pseudo magnetic field. Hence, the physics of a Chern band with Chern number $C=+1$ can be viewed as a Landau level in a magnetic field Morales-Dur¨¢n et al. (2023); Paul et al. (2023). More precisely, an electron in valley-1(2) will see magnetic flux ±$\pm\Phi_{0}=2\pi$ in each moiré unit cell. If one performs the standard flux attachment to electrons (independently for each valley), it is most natural to construct the conjugate composite Fermi liquid state discussed recently Zhang (2024a); Myerson-Jain et al. (2023); Shi et al. (2024), which is a compressible state in the bulk. However, there is another alternative flux attachment which naturally leads to an incompressible FTI.
@@
indicating the desired spin-Hall conductance $\sigma^{\mathrm{sh}}=1/2$. Also, there are no “diagonal” Hall responses for the electric charge and $S_{z}$ quantum numbers separately, which is consistent with time-reversal symmetry.
@@
The mutual Chern-Simons theory Eq. (1) implies that the mFTI is actually a $\mathbb{Z}_{4}$ TO “enriched” by two $\mathrm{U}(1)$ symmetries and time-reversal symmetry ${\cal T}$, i.e. it is a type of symmetry-enriched topological orders. As we will see below, the anyon data of this $\mathbb{Z}_{4}$ TO is equivalent to that of the $\mathbb{Z}_{4}$ toric codeKitaev (2003) or a $\mathbb{Z}_{4}$ gauge field in 2+1D. In our context, the novel aspects of the $\mathbb{Z}_{4}$ TO lie in its symmetry enrichment and the fact that it is the unique minimal FTI compatible with the symmetry and transport requirements. Throughout this work, we assume that ${\cal T}^{2}=-1$ when acting on an electron or a hole, i.e. electrons/holes are Kramers doublets. Within the $\mathbb{Z}_{4}$ TO, the “minimally-charged” anyon $\mathbf{e}$ carries physical electric charge $e/2$, and the “minimal-spin” anyon $\mathbf{m}$ carries a physical spin quantum number $S_{z}=\hbar/4$. Both of these anyons are self-bosons, but they see each other as a $\Phi_{0}/4$ flux, i.e. they have mutual statistical angle $\pi/2$. The $\mathbf{e}\mathbf{m}$ bound state carries an electric charge and a spin quantum number $(e/2,\hbar/4)$, and it is an anyon with self-statistical angle $\pi/2$, i.e. its topological spin is $\theta_{\bf em}=e^{\mathrm{i}\pi/2}=\mathrm{i}$. This $\mathbb{Z}_{4}$ TO also preserves the time-reversal symmetry ${\cal T}$, which keeps the anyon $\mathbf{e}$ invariant and maps the anyon $\mathbf{m}$ to its anti-particle $\mathbf{m}^{3}$. Curiously, $\mathbf{m}^{2}$ is a Kramers doublet under time reversal with ${\cal T}^{2}=-1$. This result can be seen using both the edge theory in Sec. III and the general arguments of Ref. Levin and Stern, 2009, which we will elaborate more on in Sec. IV.
@@
This $\mathbb{Z}_{4}$ TO have $16$ anyons modulo the electrons (and holes), which can be represented as composite states $\mathbf{e}^{p}\mathbf{m}^{q}$ with $p,q=0,\cdots 3$. However, since this $\mathbb{Z}_{4}$ TO is constructed with bosonic objects $b^{c}$ and $b^{s}$, the background gapped electrons and holes were not yet taken into account. Hence, there exist another 16 anyons, which are bound states of $\mathbf{e}^{p}\mathbf{m}^{q}$ and the electron. Note that the bound state of $\mathbf{e}^{p}\mathbf{m}^{q}$ and the electron has a different topological spin from $\mathbf{e}^{p}\mathbf{m}^{q}$. Here, we’ve effectively included the electron as a “transparent” Abelian anyon (that braids trivially with all other anyons), which is a convenient technical choice for showing the minimality of this $\mathbb{Z}_{4}$ TO later. With the electrons and holes included, the $\mathbb{Z}_{4}$ topological order has in total 32 Abelian anyons, resulting in a squared total quantum dimension ${\cal D}^{2}=32$. In Sec. IV, we will prove that ${\cal D}^{2}=32$ is the minimal squared total quantum dimension compatible with the given Hall transports associated with the electric charge and spin quantum number $S_{z}$. If one views electrons/holes as trivial anyons and identifies the anyons differed only by an electron/hole, then the mFTI effectively has a squared total quantum dimension $\tilde{{\cal D}}^{2}=16$. In the Appendix, we further prove that the mFTI proposed here is the unique mFTI compatible with charge conservation, $S_{z}$ conservation, time-reversal symmetry, and the desired Hall responses. We remark that our proof also shows that the mFTI remains to be the unique minimal TO even if we replace ${\cal T}$ by a “non-Kramers” time-reversal symmetry ${\cal T}^{\prime}$ (with ${\cal T}^{\prime 2}=1$). However, the stability of the edge state of the mFTI is different under the two types of time-reversal symmetries, as discussed in the next section.
@@ III Experimental signatures
${\cal L}_{\mathrm{edge}}=\frac{4}{4\pi}\mathrm{i}\partial_{\tau}\phi^{c}\partial_{x}\phi^{s}+\frac{4}{4\pi}\mathrm{i}\partial_{\tau}\phi^{s}\partial_{x}\phi^{c}+\cdots,$ (3)
@@
If the spin $S_{z}$ conservation is broken but time-reversal symmetry ${\cal T}$ is still preserved, an extra term $\cos(8\phi_{s})$ is allowed in the edge theory Eq. 3. When this term is relevant, it will lead to two-fold degenerate ground states at the edge, which are characterized by nonzero expectation values of gauge invariant and time-reversal odd operator $\cos(4\phi_{s})$, or $\sin(4\phi_{s})$. Hence, the edge theory remains gapless unless there is either spontaneous or explicit time-reversal symmetry breaking (assuming charge is conserved). A different way to understand the stability of the gapless edge using only the bulk data will be given below. We comment that the ${\cal T}^{2}=-1$ is important for the stability of the gapless edge (in the absence of the $S_{z}$ conservation). A non-Kramers time-reversal symmetry $\cal T^{\prime}$ would allow terms like $\cos(4\phi_{s})$ or $\sin(4\phi_{s})$ to gap out the edge without breaking $\cal T^{\prime}$. In the following analysis, the time-reversal symmetry always refers to the Kramers-type time-reversal symmetry with ${\cal T}^{2}=-1$ unless specified otherwise.
@@
${\cal L}_{\mathrm{edge}}=\frac{2}{4\pi}\mathrm{i}\partial_{\tau}\phi_{1}\partial_{x}\phi_{1}-\frac{2}{4\pi}\mathrm{i}\partial_{\tau}\phi_{2}\partial_{x}\phi_{2}+V_{IJ}\partial_{x}\phi_{I}\partial_{x}\phi_{J}.$ (4)
The density $\partial_{x}\phi_{1}$ (or $\partial_{x}\phi_{2}$) carries charge $e$ (or $-e$) and spin /2Planck-constant-over-$\hbar/2$. We have added the kinetic energy with the velocity matrix $V_{IJ}$ in Eq. (4) which arises from the non-topological part of the system. If $V_{IJ}$ is proportional to the identity matrix, the left and right counter-propagating modes do not interact with each other, and this edge theory would lead to a fractionally quantized $1d$ conductance per edge:
@@
The smallest electric charge carried by the anyons of the mFTI is $e^{\ast}=e/2$. The odd value of $\sigma^{\rm sh}/e^{*}=1$ in the mFTI guarantees the gaplessness of the edge whenever time-reversal symmetry and charge conservation are present (even without the $S_{z}$ conservation).Levin and Stern (2009) In contrast, using similar arguments as Ref. Levin and Stern, 2009, one can show that, given the spin-Hall conductance $\sigma^{\rm sh}=1/2$ (and the vanishing “diagonal” Hall responses associated with the charge and spin quantum numbers), ANY TO that is a product of two decoupled TOs from the two valleys must have excitations with fractional electric charges equal to (or smaller) than $e/4$. As a concrete example, the product of a Pfaffian state in one valley and its time-reversed partner in the other valley admits anyons with $e/4$ electric charge Moore and Read (1991); Greiter et al. (1991); Wen (1993).
Based on the discussion above, we propose the following experiments that can distinguish the mFTI from other candidates, especially the product of decoupled topological states from the two valleys:
  • Measurement of smallest charge excitations:
    The smallest fractional charge in topological states can be measured using multiple methods including the shot noise at quantum point contactsSaminadayar et al. (1997, 1997), local capacitance probe of localized charge states Martin et al. (2004), and Coulomb oscillations in antidots Kou et al. (2012). The smallest fractional charge in mFTI is $e/2$, while all the candidates formed by the products of valley-decoupled topological states must have charge-$e/4$ excitations. Therefore, the measurement of the smallest charge can distinguish the mFTI from other product topological states.
  • Anyon interferometry:
    Another characteristic feature of the mFTI is the braiding statistics $\theta_{\bf em}=e^{\mathrm{i}\pi/2}$ between the anyons ${\bf e}$ and ${\bf m}$. These braiding statistics, as well as the smallest electric charge, can be measured using anyon interferometry devicesKivelson (1990); de C. Chamon et al. (1997); Halperin et al. (2011); Nakamura et al. (2019, 2020, 2023); Willett et al. (2023); McClure et al. (2009); Camino et al. (2005, 2007); Ofek et al. (2010); Samuelson et al. (2024). In our mFTI, the ${\bf e}$ and ${\bf m}$ anyons carry electric charge and spin respectively. For a Fabry-Pérot interferometer with the mFTI, upper and lower edge currents surrounding the device interact with each other at two quantum point contacts (see Fig. 2 for a schematic illustration). The edge currents are carried by the ${\bf e}$ anyons, and the conductance of the device will receive correction from both the Aharonov–Bohm effect, as well as the braiding statistics between the ${\bf e}$ anyons in the edge currents and the localized ${\bf m}$ anyons in the device. The interference phase $\varphi$ of the current is given by
    $\varphi=2\pi\frac{e^{*}}{e}\frac{AB}{\Phi_{0}}+\frac{\pi}{2}N_{\bf m},$ (6)
    where $e^{*}=e/2$ is smallest charge in the mFTI. The first term corresponds to the Aharonov–Bohm effect proportional to the total external magnetic flux through the interferometer, where $A$ is the area of the interferometer, $B$ is the external out-of-plane magnetic field, and $\Phi_{0}$ is the magnetic flux quantum. The second term arises from the braiding statistics between ${\bf e}$ anyons in the edge current, and localized ${\bf m}$ anyons in the interferometer. $N_{\bf m}$ is the number of ${\bf m}$ anyons in the interferometer region. The factor $\pi/2$ comes from the braiding statistics $\theta_{\bf em}$ between ${\bf m}$ and the smallest charge-carrying anyon ${\bf e}$. The conductance of the interferometer, which is periodic in $\varphi$, can be measured as a continuous function of the area $A$ and the magnetic field $B$. One can extract $e^{*}$ from the periodicity of the conductance as a function of $AB$. $N_{\bf m}$ can change discretely through anyon tunneling events, resulting in jumps in the phase $\varphi$, as the anyon ${\bf m}$ jumps in or out of the interferometer region.
    In contrast, for all the product TO candidates, the smallest charge $e^{*}=e/4$ takes a different value from the mFTI, resulting in a different periodicity of $\theta$ as a function of $AB$. The term $N_{\bf m}\frac{\pi}{2}$ also needs to be modified according to the candidate states. For example, for the product of $\mathrm{U}(1)_{8}$ and its time-reversed partner, this term would be $\frac{\pi}{4}$ times the number of the smallest-charge anyon. For the product of the Pfaffian state and its time-reversal partner, one expects an even-odd effect caused by the fact that, as far as the interference pattern is concerned, an even number of the non-Abelian Ising anyons collectively yield the same statistical phase as an Ableian anyon, while an odd number of them produce interference associated with the non-Abelian statistics Bonderson et al. (2008); Willett et al. (2023).
Figure 2: A schematic illustration of a Fabry-Pérot interferometer with the mFTI. The upper and lower edges of the device are close to each other at two quantum point contacts, allowing the ${\bf e}$ to tunnel between them.